What must be the distance between point charge q1 = 26.0 micro C and point charge q2=47.0 micro C for the electrostatic force between them have a magnitude of 5.70 N?


1
Expert's answer
2021-06-01T10:50:25-0400
"F=k\frac{q_1q_2}{r^2}\\5.7=8.99\cdot10^{9}\frac{(26\cdot10^{-6})(47\cdot10^{-6})}{r^2}\\r=1.39\ m"